### v1

# Quantum Otto heat engine and quantum Mpemba effect in quasiperiodic systems

Abstract
In this paper, a quasiperiodic system with off-diagonal and diagonal modulations is studied. By analyzing the inverse participation ratio and the centroid position of the wave packet, the delocalization-localization phase diagrams in the equilibrium state and the non-equilibrium steady state are obtained respectively. The stability of the system’s initial state on dissipation and the dissipation-driven delocalization-localization transition (or vice versa) are discussed. The dynamic evolution behavior provides evidence for the existence of the quantum Mpemba effect in this dissipation-modulated system. A starting-line hypothesis is proposed, which can be used to explain the occurrence and absence of the quantum Mpemba effect. In addition, the thermodynamic applications of this system have also been studied. The results prove that the localized phase is beneficial to the realization of quantum heater as well. Our work has promoted the understanding of the steady-state phase transitions and dynamic behaviors of dissipation-modulated quasiperiodic systems, and has expanded the existing thermodynamic research on the application of quasiperiodic systems.
I Introduction
Anderson localization is a pivotal quantum phenomenon in condensed-matter physics, profoundly impacting our understanding of electron transport in disordered or quasiperiodic systems Anderson (1958); Aubry and André (1980); Mott (1987); Evers and Mirlin (2008). In three-dimensional systems, a metal-insulator transition arises based on scaling theory Abrahams et al. (1979), distinguishing a delocalized phase (with freely propagating electrons) from a localized phase (with spatially confined electron motion). Anderson localization research is vital for understanding particle behavior in complex disordered settings and advancing optical lattice experimental design and measurement techniquesSanchez-Palencia et al. (2007). Experimentally, Anderson localization has been observed in various platforms. For instance, in ultra-cold atomic systems, where precise control of interatomic interactions and external potentials enables the creation of tunable disorder and clear observation of the delocalization-to-localization transition Sanchez-Palencia et al. (2007); Roati et al. (2008); Billy et al. (2008); Kondov et al. (2011); Jendrzejewski et al. (2012); Semeghini et al. (2015); Delande and Orso (2014); Sanchez-Palencia (2015); Lüschen et al. (2018); An et al. (2021); Xiao et al. (2021). Moreover, photonic quasicrystals also exhibit Anderson localization. Studying Anderson localization in photonic crystals helps in grasping light transmission and fostering the development of functional optical devices Wiersma et al. (1997); Schwartz et al. (2007); Lahini et al. (2008); Wiersma (2013); Segev et al. (2013); Kondakci et al. (2015); Yu et al. (2021); Chang et al. (2025).
Under different physical mechanisms, Anderson localization takes various forms. In quasiperiodic systems, when the quasiperiodic potential potential strength exceeds a critical value, all quantum states become localized Aubry and André (1980); Roati et al. (2008). In one-dimensional quasiperiodic systems with short- (long-) range hoppings Biddle and Sarma (2010); Deng et al. (2019); Roy and Sharma (2021); Biddle et al. (2011); Liu et al. (2021a); Saha et al. (2019); Xu et al. (2021); Tang et al. (2021), and generalized quasiperiodic modulations Sarma et al. (1988); Ganeshan et al. (2015); Yao et al. (2019); Liu and Guo (2018); Wang et al. (2020); Liu et al. (2022); Zhou et al. (2023); Liu et al. (2020a, 2021b); Gonçalves et al. (2022, 2023a, 2023b, 2024a, 2024b); Varma et al. (2016), Anderson localization occurs only as specific energy levels, separated from delocalized levels by mobility edges. This gives the system an intermediate-phase, neither fully delocalized nor fully localized. Recently, the hidden self duality in the system with quasipriodic modulations is discovered Hu et al. (2025a), which further advances the understanding about the Anderson localization and mobility edges Li et al. (2025); Zhou et al. (2025).
Regarding to the quasiperiodic systems, the researches on mobility edges and intermediate phase have enabled applications such as energy current rectification Balachandran et al. (2019); Madhumita and Maiti (2019, 2017) and superradiance light sources Yin et al. (2020). Recently, the attention on the applications of the quasiperiodic systems are being paid to the finite-temperature aspect. It has been found that the quasiperiodic systems with mobility edges interacting with external heat baths can be used to engineer quantum heat engine Chiaracane et al. (2020); Xu and Cheng (2025). When off-diagonal quasiperiodic systems (without mobility edges) are adopted as the working medium in quantum Otto cycles Xu and Cheng (2025), a more diverse set of working modes has been identified-with the quantum heat engine being one such mode, alongside quantum accelerator and quantum heater. These findings motivate us to study when both diagonal and off-diagonal quasi-periodic modulation are introduced simultaneously and this system is used as the working medium of the quantum Otto cycle, what are the similarities and differences between the working mode of this cycle and the result when only off-diagonal modulations are considered, and what is the connection with the localized phase transition at zero temperature?
Besides the thermodynamical applications, in recent years, the burgeoning field of non-Hermitian physics Longhi (2019); Weidemann et al. (2022); Zheng et al. (2024); Liu and Chen (2024); Xu et al. (2022); Liu et al. (2020b); Purkayastha et al. (2021) and the accurate manipulation of quantum coherence in experiments Lin et al. (2022); Liang et al. (2022); Zhang et al. (2023); Zhao et al. (2025) have spurred intense interest in dissipative open quantum systems. Dissipation proves pivotal in inducing both localized and delocalized states, offering a crucial mechanism to decipher electron transport in disordered and homogeneous materials Gopalakrishnan et al. (2017); Rath and Mintert (2020); D. S. Bhakuni and Lev (2023); Longhi (2024a); Liu et al. (2024); Peng et al. (2024a). It can disrupt localization and boost transport Gopalakrishnan et al. (2017); Rath and Mintert (2020); D. S. Bhakuni and Lev (2023), induce mobility edges in systems that previously lacked delocalized-localized transitions Longhi (2024a), and drive quasiperiodic systems into specific delocalized or localized states Liu et al. (2024); Peng et al. (2024b); Hu et al. (2025b), while also modulating the topological properties of topologically trivial and nontrivial insulators Peng et al. (2024a); Feng et al. (2025), and the initial property dependence of the steady state has also been discovered Feng et al. (2025). This motivate us to study the delocalization to localization transformation of the steady state of systems with diagonal and off-diagonal quasi-periodic modulation under dissipation, and discuss the connection between the localization properties of the steady state and the initial properties of the system.
With respect to the dissipation induced delocalization-localization transition, attention are more paid on the the properties of the non-equilibrium steady states. However, the issue that the relaxation rate of systems whose initial state is further from the equilibrium state is actually faster than that of systems whose initial state is closer to the equilibrium state or steady state, namely the quantum Mpempa effect which is a generation of the Mpemba effect at the quantum domain Mpemba and Osborne (1969); Kell (1969); Jeng (2006); Bechhoefer et al. (2021); Teza et al. (2025), remains to be studied. Regarding to the quantum Mpemba effect, it has been theoretically predicted across a diverse range of systems-including quantum dots Chatterjee et al. (2023); Wang and Wang (2024); Graf et al. (2025); Strachan et al. (2024), spin or bosonic systems Longhi (2024b); Nava and Fabrizio (2019); Caceffo et al. (2024); Moroder et al. (2024); Bhore et al. (2025); Westhoff et al. (2025), multi-level systems Carollo et al. (2021); Manikandan (2021); Ivander et al. (2023); Chatterjee et al. (2024), as well as other platforms Chétrite et al. (2021); Holtzman and Raz (2022); Dong et al. (2025); Di Giulio et al. (2025). Experimentally, this effect has also been observed in specific setups: trapped-ion quantum simulators Joshi et al. (2024) and single trapped ions Aharony Shapira et al. (2024); Zhang et al. (2025). Motivated by these outstanding works, we want to know whether the quantum Mpemba effect will occur in this open system with diagonal and off-diagonal quasiperiodic modulations as well. A previous study has shown that if the initial state is localized, the quantum Mpemba effect will occur Dong et al. (2025). Therefore, we want to study whether the quantum Mpemba effect will occur if the initial state is not a localized state as well. Furthermore, we also aim to provide a unified explanation for the appearance and absence of the quantum Mpemba effect.
This work is organized as follows. Section II studies the model with diagonal and off-diagonal modulations and gives out the localization phase diagram. Section III studies the thermodynamical applications of the quantum Otto cycle with the quasiperiodic system being the working medium. Section IV studies the delocalization-localization transition of the steady state after considering the bond dissipations. Section V reveals the phenomenon of quantum Mpemba effect and puts forward a hypothesis to explain the cause of this phenomenon. A summary is presented in Sec. VI.
Figure 1: (Color Online) The phase diagram illustrates how $\log_{10}(\rm MIPR)$ varies with $V/t$ and $t_{1}/t$, where the system size is set to $L=1597$. The red dashed line represents the phase boundary, which serves to distinguish the extended phase from the localized and critical phases. Meanwhile, the colorbar corresponds to the numerical values of $\log_{10}(\rm MIPR)$.
II Model and localization phase diagram
We study the quasiperiodic system with off-diagonal and diagonal quasiperiodic modulation, whose Hamiltonian is presented as
$H=\sum_{n}\left(t_{n}\hat{c}^{{\dagger}}_{n}\hat{c}_{n+1}+H.c.\right)+\sum_{n}V_{n}\hat{c}^{{\dagger}}_{n}\hat{c}_{n},$ (1)
where $t_{n}=t+t_{1}\cos{(2\pi\alpha n)}$ and $V_{n}=V\cos{(2\pi\alpha n)}$ ($t$ is the unit of energy, $n$ is the site index, and $\alpha=\frac{\sqrt{5}-1}{2}$ is the incommensurate parameter which is the cause of the quasiperiodic modulations.).
Figure 2: (Color Online)(a) $\log_{10}(\rm MIPR)$ as a function of $r$ with $\theta=0$, $\pi/4$, $\pi/3$, and $5\pi/6$. (b) $\overline{D_{min}}$ as a function of the inverse Fibonacci index $1/m$ under different $(r,\theta)$.
The localization phase diagram can be determined from the inverse participation ratio (IPR). Given a normalized wave function $\psi_{j}=\sum^{L}_{n=1}\phi_{j}(n)\hat{c}^{{\dagger}}_{n}|0\rangle$ (Here $j$ is the index of wave function), the corresponding ${\rm IPR}_{j}$ is
${\rm IPR}_{j}=\sum^{L}_{n=1}|\phi_{j}(n)|^{4}.$ (2)
It is known that the the $\rm IPR$} of the extended state, critical state, and localized state respectively possess the properties of ${\rm IPR}\approx 0$, $0<{\rm IPR}<1$, and ${\rm IPR}\approx 1$ Biddle and Sarma (2010). Under the given parameters, the extended phase, critical phase and localized phase can be further characterized by the average value of $\rm IPR$} of all wave functions, namely the average inverse participation ratio (MIPR), which is denoted as ${\rm MIPR}=\sum^{L}_{j=1}{\rm IPR}_{j}/L$. Taking system size $L=1597$, the localization phase diagram in $V$-$t_{1}$ parameter space is plotted in Fig. 1, where the color bar denotes $\log_{10}(\rm MIPR)$. In fact, for the $V=0$ case, the earlier studies have proved that the system exits an extended-critical phase transition Liu et al. (2016). We can see from Fig. 1 that under $V=0$, $\log_{10}(\rm MIPR)$ for the extended phase approaching 3-3 and $\log_{10}(\rm MIPR)$ for the critical phase approaches 1-1. Based on this, we can infer that in the case of a finite $V$, the quarter elliptic region with $\log_{10}(\rm MIPR)$ close to 3-3 should be the extended phase. On its outside, $\log_{10}(\rm MIPR)$ is greater than 1-1, so it should be a localized phase.
Compared with the non-Hermitian situation Chen et al. (2022), the recovery of system’s Hermiticity expands the area of the extended phase. To locate the phase boundary of the extended-localized transition, we shall introduce the transformation $r=\sqrt{\frac{V^{2}}{4}+t^{2}_{1}}$ with $\theta={\rm Arctan}(V/t_{1})\in\left[0,2\pi\right]$. Under $L=1597$, the $\log_{10}(\rm MIPR)$ versus $r$ with different $\theta$ are plotted in Fig. 2(a). Intuitively, we can see that for different $\theta$, the $\log_{10}(\rm MIPR)$-$r$ curves characterizing the extended-localized (critical) transition all present a jump at the critical value $r_{c}/t=1$, namely the phase boundary $\sqrt{\frac{V^{2}}{4}+t^{2}_{1}}=t$, which has been plotted as the red dashed line in Fig. 1.
To further verify the above mentioned conclusions we have drawn, we introduce the fraction dimension $D$. Consider a system where the system size $L$ equals the $m$-th Fibnacci number $F_{m}$ and the incommensurate parameter $\alpha$ is replaced by $\alpha=F_{m-1}/F_{m}$. At a specific lattice site $n$, the fraction dimension $D_{n}$ can be derived from the formula
$p_{n}=F_{m}^{-D_{n}}.$ (3)
Here, $p_{n}$ represents the probability density. This equation reveals that $D_{n}$ acts as a scaling index. For an extended state, since $p_{n}\sim 1/F_{m}$, we have $D_{n}\sim 1$. In the case of a localized state, $D_{n}\sim 1$ at the sites where the particle occupies, while $D_{n}\rightarrow\infty$ at unoccupied sites. For a critical state, $D_{n}$ lies within a finite interval $\left[D_{min},D_{max}\right]$. Thus, the minimum value $D_{min}$ directly reflects the nature of a given wave function $\psi_{j}$. Specifically, $D_{min}\rightarrow 0$ indicates a localized state, $0<D_{min}<$ points to a critical state, and $D_{min}\rightarrow 1$ signifies an extended state. For generality, we use the average of $D_{min}$ over all states, defined as $\overline{D_{min}}=\sum^{L}_{j=1}D^{j}_{min}/L$. In the limit of $1/m\rightarrow 0$ (extrapolation limit), this average helps distinguish different phases. We select representative parameter points in various phases to compute $\overline{D_{min}}$. As depicted in Fig. 2(b), we observe that the relevant $D_{min}$ approaches $1$ at the parameter sites $\left(r,\theta\right)=\left(0.5t,\frac{\pi}{4}\right)$ and $\left(r,\theta\right)=\left(0.8t,\frac{\pi}{3}\right)$. This observations confirm that the system resides in the extended phase under these conditions. As anticipated, the corresponding $D_{min}$ values fall within the interval $\left(0,1\right)$ in the thermodynamic limit when $\left(r,\theta\right)=\left(1.5t,0\right)$ and $\left(r,\theta\right)=\left(2t,0\right)$. These results clearly reveal the systems’s critical-phase characteristics in such parameter regimes. When $\left(r,\theta\right)=\left(1.5t,\frac{\pi}{3}\right)$ and $\left(r,\theta\right)=\left(2t,\frac{\pi}{6}\right)$, the associated $D_{min}$ extrapolates to 0. This outcome identifies that the system is in the localized phase at these parameter points.
III Thermodynamical applications
In recent years, some researches have been carried out on the thermodynamical applications of quasi-periodic systems Chiaracane et al. (2020); Xu and Cheng (2025), leading to the identification of their applications in aspects such as quantum heat engine, quantum heater, and quantum accelerator. Specifically, it has been demonstrated that the extended phase plays a role in sustaining the operational mode of heat engine, whereas the critical phase is conducive to maintaining the working modes of quantum heater and quantum accelerator. Within this section, we focus on investigating the thermodynamic applications of this generalized Aubry-André (AA) model with both diagonal and off-diagonal quasiperiodic modulations. Our primary objectives include verifying whether the extended phase of this model still contributes to preserving the working mode of quantum heat engines, as well as determining which specific working mode the localized phase favors.
Figure 3: (Color Online) Schematic illustration of the four-stroke quantum heat cycle process. $T_{h}$ and $T_{c}$ stand for the high-temperature and low-temperature heat sources, respectively. The working medium employed is the generalized AA model with diagonal and off-diagonal quasiperiodic modulations. $r^{i}$ and $r^{f}$ represent the systematic parameters of the corresponding Hamiltonians. The system size is $L=610$.
We leverage this quasiperiodic system as the working medium to construct a quantum heat cycle process. The schematic of this cycle is shown in Fig. 3. This quantum heat cycle can be regarded as the quantum counterpart Quan et al. (2007); Niskanen et al. (2007); Abah et al. (2012); Deng et al. (2013); del Campo et al. (2014); Roßnagel et al. (2016); Zheng et al. (2016); Kosloff and Rezek (2017); Deffner and Campbell (2019); Campisi and Fazio (2016); Buffoni et al. (2019); Solfanelli et al. (2020); Zhang et al. (2022) of the classical Otto cycle. The first (④ $\rightarrow$ ①) and third (② $\rightarrow$ ③) strokes occur under thermal contact with high-temperature ($T_{c}$) heat reservoirs. These strokes proceed without external driving or particle exchange. The second (① $\rightarrow$ ②) and fourth (③ $\rightarrow$ ④) strokes are adiabatic from a thermodynamic viewpoint (meaning thermal isolation), yet they might not be strictly adiabatic in the quantum-mechanical sense, since quantum transitions can occur during the evolution [82].
In the first stroke, the working medium, characterized by the Hamiltonian $H(r^{i})$, will eventually reach thermal equilibrium and relax into a Gibbs state. The density matrix $\rho_{1}$ of this state is given by $\rho_{1}=\frac{e^{-\beta_{h}H(r^{i})}}{Z_{1}}$, where $\beta_{h}=\frac{1}{k_{b}T_{h}}$ ($k_{b}$ is Boltzmann constant) and $Z_{1}={\rm Tr}\left[e^{-\beta_{h}H(r^{i})}\right]$ is the partition function. Thus, at thermal equilibrium, the system energy is $E_{1}={\rm Tr}\left[\rho_{1}H(r^{i})\right]$. During the second stroke, the Hamiltonian parameter is switched from $r^{i}$ to $r_{f}$. In this stroke, only work is done, and no heat exchange takes place. The Gibbs state $\rho_{2}$ of this stroke remains unchanged, i.e, $\rho_{2}=\rho_{1}$, but the energy $E_{2}$ becomes $E_{2}={\rm Tr}\left[\rho_{2}H(r)\right]$. In the third stroke, the medium with $r=r^{f}$ comes into contact with a heat source having $\beta_{c}=\frac{1}{k_{b}T_{c}}$. As a result, the Gibbs state $\rho_{3}=\frac{e^{-\beta_{c}H(r^{f})}}{Z_{2}}$, with the partition function $Z_{2}={\rm Tr}\left[e^{-\beta_{c}H(r^{f})}\right]$. Then, the energy of the medium is $E_{3}={\rm Tr}\left[\rho_{3}H(r^{f})\right]$. The fourth stroke can be regarded as the thermal annealing process. In this process, the hopping parameter in the medium is changed back from $r^{f}$ to $r^{i}$. However, the Gibbs state $\rho_{4}$ stays as $\rho_{4}=\rho_{3}$. Hence, the energy of the medium becomes $E_{4}={\rm Tr}\left[\rho_{4}H(r^{i})\right]$.
Figure 4: (Color Online) Working modes of the four-stroke cycle. (a) $r^{f}=0.5t$ and $\theta=\pi/6$. (b) $r^{f}=0.5t$ and $\theta=\pi/4$. (c) $r^{f}=0.9t$ and $\theta=\pi/6$. (d) $r^{f}=0.9t$ and $\theta=\pi/4$. (e) $r^{f}=1.1t$ and $\theta=\pi/6$. (f) $r^{f}=1.1t$ and $\theta=\pi/4$. The blue regions represent the Heat engine. The yellow regions denote the Accelerator and the brown regions denote the Heater. Other parameters are $L=610$ and $k_{b}T_{c}=0.1t$.
After the working medium undergoes this complete cycle, we can determine the heat absorbed $Q_{h}$ ($Q_{h}=E_{1}-E_{4}$) from the $T_{h}$ source, the heat released $Q_{c}$ ($Q_{c}=E_{3}-E_{2}$) to the $T_{c}$ source, and the net work $W$ done by the working medium, where $W=Q_{h}+Q_{c}$. Importantly, the heat cycle process abides by the Clausius inequality, a cornerstone of thermodynamics. Depending on the values of $Q_{h}$, $Q_{c}$, and $W$, the engine using the extended-localized (critical) quasiperiodic system as the working medium exhibits different modes Buffoni et al. (2019); Solfanelli et al. (2020):
  1. 1.
    Heat engine: $Q_{h}>0$, $Q_{c}<0$, and $W>0$;
  2. 2.
    Refrigerator: $Q_{h}<0$, $Q_{c}>0$, and $W<0$;
  3. 3.
    Heater: $Q_{h}<0$, $Q_{c}<0$, and $W<0$;
  4. 4.
    Accelerator: $Q_{h}>0$, $Q_{c}<0$, and $W<0$.
For a system with size $L=610$ and $k_{b}T_{c}=0.1t$, after analyzing the values of $Q_{h}$, $Q_{c}$ and $W$, the corresponding working modes of the four-stroke cycle under different $r^{f}$ and $\theta$ are plotted in Figs. 4(a)-4(f), respectively. It can be observed that this cycle exhibits a diverse range of working modes. The blue regions represent the Heat engine mode. The yellow regions correspond to the Accelerator mode, while the brown regions indicate the Heater working mode. It is noted from Figs. 4(a)-4(d) that when $r^{f}$ and $r^{i}$ are below the critical value $r_{c}$, the Heat engine mode appears. This suggests that similar to the results of the critical-extended quasi-periodic system Xu and Cheng (2025), for the quasi-periodic system under a finite $V$, its extended phase is favorable for the design of a quantum heat engine as well. As $r^{f}$ increases and surpasses the critical values $r_{c}=t$, the system entries the localized phase. From Figs. 4(e) and 4(f), we can see that regions representing the Heater mode expand. This indicates that the localized phase is more conducive to the realization of a quantum heater. Additionally, it is clear that there are extensive parameter ranges corresponding to the Accelerator mode, regardless of whether the system is in the extended phase or the localized phase. Combining with previous research on extended-critical quasiperiodic system Xu and Cheng (2025), we draw a conclusion that all the extended, critical and localized phases facilitate the realization of a quantum accelerator, and both the localized and critical phases are conducive to the realization of quantum heater.
Moreover, the results shown in Fig. 4 offer strategies for regulating the working mode of the four-stroke cycle. The transition between different working modes can be achieved by tuning $r^{f}$, $r^{i}$, and $k_{b}T_{h}$. For example, when $r^{f}$ is much smaller than $r_{c}$ ( refer to Figs. 4(a) and 4(b)), or vice versa, there are two working modes, i.e., Accelerator and Heat engine. By tuning $r^{i}$ and $k_{b}T_{h}$, the Otto cycle can switch between the two working modes. When $r^{f}$ is close to but still less than $r^{c}$ and $k_{b}T_{h}$ is small (refer to Figs. 4(c) and 4(d)), there are three working modes. Thus, the cycle’s working mode will change from Heater to Accelerator and then to Heat engine when we gradually increase $r^{i}$. When $r^{f}$ is larger than $r_{c}$, the cycle can be toggled between Heater and Accelerator modes by tuning $r^{i}$ or $k_{b}T_{h}$.
IV Dissipation driven delocalization-localization transition
In the preceding section, we conducted a systematic investigation into the thermodynamic characteristics of this quasi-periodic system under thermal equilibrium conditions. Our findings revealed two key insights: firstly, the system’s thermodynamic behaviors are tied to its initial properties; secondly, the system holds potential application value for developing quantum heat engine, heater, and accelerator. We also analyzed approaches to regulate these three working modes. In addition to the thermodynamic aspect, recent research efforts have yielded conflicting perspectives regarding the steady-state delocalization-localization transition under dissipative modulation Longhi (2024a); Liu et al. (2024); Peng et al. (2024b); Hu et al. (2025b). By employing the dissipation, one can precisely control the delocalization and localization properties of the system under non-equilibrium steady states. Therefore, it motivates us to study the delocalization and localization properties of the steady state after introducing the dissipation.
Besides, some studies suggest this transition is independent of the system’s initial properties. Alternative viewpoints argue that the system’s initial attributes can influence the transition. Different initial property may either shorten or extend the threshold required for the delocalization-localization shift to occur. We aim to study whether the dependence of system’s initial property will also occur in this system with diagonal and off-diagonal quasiperiodic modulations and to study how different initial properties affect the delocalization-localization transition of the steady state as well.
Upon introducing this bond dissipation Lindblad operator $L_{n}=c^{{\dagger}}_{n}c_{n+\ell}$, generated following the ones in Ref.Yusipov et al. (2017); Vershinina, O. S. et al. (2017); Yusipov et al. (2018); Vakulchyk et al. (2018); Jiang et al. (2024); Yang et al. (2024); Liu et al. (2024), then the time-evolution of the density matrix $\rho$ is dictated by the following Lindblad master equation
$\dot{\rho}=-\left[H,\rho\right]+\sum_{n}\gamma_{n}\mathcal{D}\left[L_{n}\right]\rho,$ (4)
Here, the dissipator $\mathcal{D}\left[L_{n}\right]$ is defined as $\mathcal{D}\left[L_{n}\right]=L_{n}\rho L^{{\dagger}}_{n}-\{L^{{\dagger}}_{n}L_{n},\rho\}/2$, and $\gamma_{n}=\gamma\cos({2\pi\alpha n})$ represents the tunably quasiperidoic bond dissipation strength with the incommensurate parameter $\alpha=\frac{\sqrt{5}-1}{2}$. The rationale for our selection of the quasiperiodic form for the key dissipation operator stems from the consideration of the following aspect: the initial key dissipation operator is made up of four terms, with two being dephasing terms Longhi (2024a) and the remaining two being hopping terms. As indicated in Ref. Liu et al. (2024), the dephasing effect fails to induce the delocalization-localization transition of steady states. It has been theoretically and experimentally confirmed that quasiperiodic modulation is capable of causing localization Roati et al. (2008). Moreover, this bond dissipation operator can be experimentally realized by introducing the auxiliary lattice Diehl et al. (2008); Kraus et al. (2008); Diehl et al. (2010, 2011); Bardyn et al. (2013). Owing to these factors, we utilized such a dissipative operator. This evolution equation can also be equivalently expressed as
$\rho(\tau)=e^{\mathcal{L}\tau}\rho(0),$ (5)
which encapsulates all the system’s dynamical information throughout the evolution process. Given that the real components of the eigenvalues of the Lindbladian matrix $\mathcal{L}$ are non-positive, as time evolves to the long-time limit, the density matrix will ultimately relax to the zero-energy eigenstate of $\mathcal{L}$, namely the non-equilibrium steady state (NESS) $\rho_{ss}$.
Figure 5: (Color Online) Steady state phase diagram. (a) $V=0$. (b) $V=0.5t$. (c) $V=t$. (d) $V=2t$. The blue regions denote the delocalization phases. The yellow regions denote the localization phases. Other parameter is $L=100$.
With $L=100$ and $\ell=1$, the delocalization phase diagrams for $V=0$, $V=0.5t$, $V=t$, and $V=2t$ are plotted in Figs. 5(a)- 5(d), respectively. The phase diagram is obtained by analyzing the position of the centroid in a steady state. We define a steady state where the centroid position is no more than five times the lattice constant from the system boundary as a localized state; otherwise, it is a delocalized state. The reason for choosing such a five-fold distance as the criterion is based on the experimental fact that in Roati et al. (2008), the authors reproduced a one-dimensional AA model in the cold atom optical lattice experimental system and observed the Anderson localization phase transition, where the wave packet broadening of the localized state was approximately $10$ times the lattice constant Roati et al. (2008). Therefore, it is feasible to distinguish the delocalization-localization transition of the steady state through the position of the center of mass.
As can be seen from Figs. 5(a)-5(c), when $t_{1}$ is less than $0.5t$, there is a large parameter range that makes all the steady states delocalized. This indicates that under these parameters, the steady state can retain the initial delocalized property of the system. As $\gamma$ or $t_{1}$ increases, the steady state gradually changes from a delocalized state to a localized state, meaning that the properties of the steady state can be regulated by tuning $\gamma$ and $t_{1}$. Meanwhile, we observe that as $V$ gradually increases, the range of the delocalized region gradually shrinks. It is indicated that adjusting the parameter $V$ can also regulate the localization property of the steady state. For instance, when V is at a relatively small value, it can be seen from 5(a) and 5(b) that when $t_{1}>t$, there are still delocalized states within it, resulting in the reentrance delocalization phenomenon. Moreover, the value of $V$ is smaller, making the delocalization region larger. When $V$ is relatively large, it can be seen from Figs. 5(c) and 5(d) that the delocalized region disappears. In particular, we find that when $V=2t$, even if the initial system is in the localized phase, the steady state is still delocalized (as can be seen from Fig. 5(d), the delocalized state exists within a small range of $t_{1}$ and $\gamma$).
V quantum Mpemba effect and starting-line hypothesis
Regarding to the delocalization-localization transition in open quasiperiodic system, the focus on this topic mainly lies in the localization properties of steady states. For an anomalous dynamic process: the relaxation rate of a system whose initial state is further from the equilibrium state is actually faster than that of a system whose initial state is closer to the equilibrium state, that is, the quantum Mpemba effect, remains to be studied and discussed.
Figure 6: (Color Online) Trace distance $d_{\rm Tr}$ versus evolution time $\tau$. (a) $V=0$, $t_{1}=1.2t$, and $\gamma=0.5t$. (b) $V=0$, $t_{1}=1.7t$, and $\gamma=0.8t$. (c) $V=0.2t$, $t_{1}=1.2t$, and $\gamma=0.2t$. (d) $V=0.2t$, $t_{1}=1.5t$, and $\gamma=0.2t$. (e) $V=0.2t$, $t_{1}=0.8$, $\gamma=0.2t$. (f) $V=0.5t$, $t_{1}=0.5$, $\gamma=0.2t$. The blue regions denote the delocalization phases. The yellow regions denote the localization phases. Other parameter is $L=100$.
We employ the density matrix $\rho(\tau)$ that describe the evolution of the system over time and the steady state density matrix $\rho_{ss}$ to calculate the standard distance between the matrices, this is, the trace distance, to characterize the quantum Mpemba effect. The trace distance $d_{\rm Tr}$ is defined as
$d_{\rm Tr}=\frac{1}{2}{\rm Tr}\sqrt{T^{{\dagger}}T},$ (6)
where $T=\rho(\tau)-\rho_{ss}$. Under Lindbladian dynamics, this quantity displays monotonic behavior over time. Notably, it has already been utilized as an experimental probe to investigate both the strong quantum Mpemba effect and its inverse counterpart Aharony Shapira et al. (2024); Zhang et al. (2025).
Prior to investigating the quantum Mpemba effect, we first introduce the ‘starting-line hypothesis’. This hypothesis posits the following: for an initial state whose energy is closest to the steady-state energy, if the distance between its center of mass and the steady-state center of mass is larger than that of other initial states with energies further from the steady-state energy, then the quantum Mpemba effect will arise. Before verifying this hypothesis, we can draw a simple inference regarding which initial states are capable of triggering the quantum Mpemba effect. For quantum states with extended properties, the position of their center of mass is consistently situated at the center of the system. In contrast, for quantum states with localized and critical properties, the positions of their centroids are not fixed at the center of system; instead, they can be distributed at other locations within the system. As a result, when quantum states with localized and critical properties are chosen as initial states, they will encounter the ‘starting-line discrepancy’ a condition that gives rise to the quantum Mpemba effect. On the other hand, for initial states with extended properties, their centroids all align at the same ‘starting-line’ (i.e., the center of system), meaning the quantum Mpemba effect will not occur.
We take three different types of system eigenstates as the initial states and calculated the evolution of trace distance over time, as shown in Figures 6(a)-6(f). The ordinal number in the caption indicates the ordinal number of the initial state in the eigenstates of $H$. Specifically, the initial state corresponding to the red curve is the eigenstate of the system that is closest to the steady-state energy. Figures 6(a) and 6(b) show the results of the local state as the initial state. The centroid positions corresponding to the initial state (rounded off, the same below) are respectively: $n=\{60,49,39\}$ and $n=\{60,49,41\}$. Figures 6(c) and 6(d) show the results of delocalized states with critical properties as the initial states, where the corresponding centroid positions of the initial states are respectively: $n=\{63,61,30,83,49\}$ and $n=\{66,63,15,59,71\}$. Figures 6(e) and 6(f) present delocalized states with critical properties as the results of initial states, where the corresponding centroid positions of the initial states are respectively: all $n=50$ and all $n=51$. We can observe that for the initial states with extended properties, the dynamical evolution behavior does not exhibit the quantum Mpemba effect, while for the initial states with critical and localized properties, the dynamical evolution behavior will exhibit the quantum Mpemba effect. From the perspective of the centroid position of the initial state, when the Mpemba effect occurs, the wave packet of the initial state is relatively closer to the boundary of the system, which confirms the starting-line hypothesis we proposed.
VI Summary
For one-dimensional chain models where both hoppings and onsite potential are quasiperiodically modulated, we have analyzed their delocalization-localization transitions under equilibrium and non-equilibrium conditions, thermodynamic applications, and dynamic evolution behaviors. We numerically calculated and obtained the delocalization-localization phase diagram of the model, and verified it through MIPR and fractal dimensions. The results show that the delocalized phase with extended feature and and the localized phase can be separated by an ellipse curve. By using the model as the working medium of the quantum Otto cycle, we analyzed the underlying working modes of the cycle under multiple sets of typical parameters. The results show that when the system is in the delocalized phase with extended feature, it is conducive to the realization of the quantum heat engine, while being in the localized phase is conducive to the realization of the quantum heater. Furthermore, regardless of the phase region where the system is located, the working parameter space of the accelerator mode is always the largest, while that of the refrigerator mode never exists. After calculating the delocalization-localization phase diagram of the non-equilibrium steady state, we discussed the influence of quasiperiodic bond dissipation on the initial state evolution of the model and found that under moderate $V$ and $t_{1}$, the steady state of the system can largely maintain the initial delocalization property of the system. By adjusting $\gamma$ and $t_{1}$ or by adjusting the parameter $V$, the properties of the steady state can be regulated, gradually changing it from a delocalized state to a localized state. When the value of $V$ is much smaller, even if the initial state is localized, the steady state can also be delocalized. We analyzed the non-equilibrium dynamic evolution behavior and the existence of the quantum Mpemba effect using the trace distance. The results told that the quantum states with localized and critical properties as initial states are necessary conditions for the occurrence of the quantum Mpemba effect. These results can be explained by the starting-line hypothesis we have proposed. As for whether this hypothesis still holds true for other open quasi-periodic systems, this is an open question that remains to be tested.
This research is supported by Zhejiang Provincial Natural Science Foundation of China under Grant No. LQN25A040012, the National Natural Science Foundation of China under Grant No. 12174346, and the start-up fund from Xingzhi College, Zhejiang Normal University.

### v3

@@
-Quantum Otto heat engine and quantum Mpemba effect in quasiperiodic systems
+Quantum Mpemba effect in quasiperiodic systems
 Abstract
-In this paper, a quasiperiodic system with off-diagonal and diagonal modulations is studied.
-By analyzing the inverse participation ratio and the centroid position of the wave packet, the delocalization-localization phase diagrams in the equilibrium state and the non-equilibrium steady state are obtained respectively.
-The stability of the system’s initial state on dissipation and the dissipation-driven delocalization-localization transition (or vice versa) are discussed.
-The dynamic evolution behavior provides evidence for the existence of the quantum Mpemba effect in this dissipation-modulated system.
-A starting-line hypothesis is proposed, which can be used to explain the occurrence and absence of the quantum Mpemba effect.
-In addition, the thermodynamic applications of this system have also been studied.
-The results prove that the localized phase is beneficial to the realization of quantum heater as well.
-Our work has promoted the understanding of the steady-state phase transitions and dynamic behaviors of dissipation-modulated quasiperiodic systems, and has expanded the existing thermodynamic research on the application of quasiperiodic systems.
+We study a one-dimensional quasiperiodic tight-binding model with simultaneous off-diagonal (hopping) and diagonal (onsite) modulations.
+Using the inverse participation ratio and the wave-packet centroid, we construct localization–delocalization phase diagrams for both equilibrium and nonequilibrium steady states.
+We analyze the robustness of initial-state properties under dissipation and characterize dissipation-induced localization–delocalization transitions (and their reversals) in detail.
+Trace-distance dynamics provide evidence for a quantum Mpemba effect: states prepared farther from the steady state can relax faster than states initialized closer to it.
+We propose a starting-line hypothesis that explains the presence or absence of this effect across parameter regimes.
+These results advance the understanding of steady-state phase transitions and relaxation dynamics in dissipatively driven quasiperiodic systems.
 I Introduction
@@
 Under different physical mechanisms, Anderson localization takes various forms.
-In quasiperiodic systems, when the quasiperiodic potential potential strength exceeds a critical value, all quantum states become localized Aubry and André (1980); Roati et al.
+In quasiperiodic systems, when the quasiperiodic potential strength exceeds a critical value, all quantum states become localized Aubry and André (1980); Roati et al.
 (2008).
@@
 (2011); Liu et al.
-(2021a); Saha et al.
+(2021); Saha et al.
 (2019); Xu et al.
 (2021); Tang et al.
-(2021), and generalized quasiperiodic modulations Sarma et al.
+(2021) and generalized quasiperiodic modulations Sarma et al.
 (1988); Ganeshan et al.
-(2015); Yao et al.
-(2019); Liu and Guo (2018); Wang et al.
-(2020); Liu et al.
+(2015); Wang et al.
 (2020); Zhou et al.
 (2023); Liu et al.
-(2020a, 2021b); Gonçalves et al.
-(2022, 2023a, 2023b, 2024a, 2024b); Varma et al.
-(2016), Anderson localization occurs only as specific energy levels, separated from delocalized levels by mobility edges.
+(2020); Gonçalves et al.
+(2023, 2024); Varma et al.
+(2016) , Anderson localization occurs only as specific energy levels, separated from delocalized level by mobility edges.
 This gives the system an intermediate-phase, neither fully delocalized nor fully localized.
-Recently, the hidden self duality in the system with quasipriodic modulations is discovered Hu et al.
+Recently, the hidden self duality in the system with quasiperiodic modulations is discovered Hu et al.
 (2025a), which further advances the understanding about the Anderson localization and mobility edges Li et al.
@@
 (2025).
-Regarding to the quasiperiodic systems, the researches on mobility edges and intermediate phase have enabled applications such as energy current rectification Balachandran et al.
-(2019); Madhumita and Maiti (2019, 2017) and superradiance light sources Yin et al.
-(2020).
-Recently, the attention on the applications of the quasiperiodic systems are being paid to the finite-temperature aspect.
-It has been found that the quasiperiodic systems with mobility edges interacting with external heat baths can be used to engineer quantum heat engine Chiaracane et al.
-(2020); Xu and Cheng (2025).
-When off-diagonal quasiperiodic systems (without mobility edges) are adopted as the working medium in quantum Otto cycles Xu and Cheng (2025), a more diverse set of working modes has been identified-with the quantum heat engine being one such mode, alongside quantum accelerator and quantum heater.
-These findings motivate us to study when both diagonal and off-diagonal quasi-periodic modulation are introduced simultaneously and this system is used as the working medium of the quantum Otto cycle, what are the similarities and differences between the working mode of this cycle and the result when only off-diagonal modulations are considered, and what is the connection with the localized phase transition at zero temperature?
-Besides the thermodynamical applications, in recent years, the burgeoning field of non-Hermitian physics Longhi (2019); Weidemann et al.
-(2022); Zheng et al.
-(2024); Liu and Chen (2024); Xu et al.
-(2022); Liu et al.
-(2020b); Purkayastha et al.
-(2021) and the accurate manipulation of quantum coherence in experiments Lin et al.
+As a powerful framework for addressing dissipation Breuer and Petruccione (2007); Ashida et al.
+(2020), the burgeoning field of accurate manipulation of quantum coherence in experiments Lin et al.
 (2022); Liang et al.
@@
 Dissipation proves pivotal in inducing both localized and delocalized states, offering a crucial mechanism to decipher electron transport in disordered and homogeneous materials Gopalakrishnan et al.
-(2017); Rath and Mintert (2020); D.
-S.
-Bhakuni and Lev (2023); Longhi (2024a); Liu et al.
-(2024); Peng et al.
-(2024a).
+(2017); Rath and Mintert (2020); Bhakuni et al.
+(2024); Longhi (2024); Liu et al.
+(2024) .
 It can disrupt localization and boost transport Gopalakrishnan et al.
-(2017); Rath and Mintert (2020); D.
-S.
-Bhakuni and Lev (2023), induce mobility edges in systems that previously lacked delocalized-localized transitions Longhi (2024a), and drive quasiperiodic systems into specific delocalized or localized states Liu et al.
-(2024); Peng et al.
-(2024b); Hu et al.
-(2025b), while also modulating the topological properties of topologically trivial and nontrivial insulators Peng et al.
-(2024a); Feng et al.
+(2017); Rath and Mintert (2020); Bhakuni et al.
+(2024), induce mobility edges in systems that previously lacked delocalized-localized transitions Longhi (2024), and drive quasiperiodic systems into specific delocalized or localized states Liu et al.
+(2024); Hu et al.
+(2025b), while also modulating the topological properties of topologically trivial and nontrivial insulators Feng et al.
 (2025) , and the initial property dependence of the steady state has also been discovered Feng et al.
@@
 This motivate us to study the delocalization to localization transformation of the steady state of systems with diagonal and off-diagonal quasi-periodic modulation under dissipation, and discuss the connection between the localization properties of the steady state and the initial properties of the system.
-With respect to the dissipation induced delocalization-localization transition, attention are more paid on the the properties of the non-equilibrium steady states.
-However, the issue that the relaxation rate of systems whose initial state is further from the equilibrium state is actually faster than that of systems whose initial state is closer to the equilibrium state or steady state, namely the quantum Mpempa effect which is a generation of the Mpemba effect at the quantum domain Mpemba and Osborne (1969); Kell (1969); Jeng (2006); Bechhoefer et al.
+In studies of dissipation-induced delocalization–localization transitions, prior work has largely concentrated on the properties of nonequilibrium steady states.
+Indeed, the possibility that states initialized farther from the steady (equilibrium) state may relax faster than those prepared closer to it—an instance of the quantum Mpemba effect, the quantum counterpart of the classical phenomenon—has also attracted considerable attention Mpemba and Osborne (1969); Kell (1969); Jeng (2006); Bechhoefer et al.
 (2021); Teza et al.
-(2025), remains to be studied.
-Regarding to the quantum Mpemba effect, it has been theoretically predicted across a diverse range of systems-including quantum dots Chatterjee et al.
+(2026).
+Regarding to the quantum Mpemba effect, it has been theoretically predicted across a diverse range of systems, including quantum dots Chatterjee et al.
 (2023); Wang and Wang (2024); Graf et al.
 (2025); Strachan et al.
-(2024), spin or bosonic systems Longhi (2024b); Nava and Fabrizio (2019); Caceffo et al.
-(2024); Moroder et al.
+(2025), spin or bosonic systems Nava and Fabrizio (2019); Moroder et al.
 (2024); Bhore et al.
@@
 (2023); Chatterjee et al.
-(2024), as well as other platforms Chétrite et al.
-(2021); Holtzman and Raz (2022); Dong et al.
-(2025); Di Giulio et al.
+(2024), as well as other platforms Holtzman and Raz (2022); Dong et al.
+(2025); Wei et al.
+(2025); Xu et al.
+(2025); Turkeshi et al.
 (2025).
-Experimentally, this effect has also been observed in specific setups: trapped-ion quantum simulators Joshi et al.
-(2024) and single trapped ions Aharony Shapira et al.
+A theoretical framework for the origin of the Mpemba effect in closed many-body quantum systems, together with concise reviews and perspectives on its quantum variants, has been presented Rylands et al.
+(2024); Ares et al.
+(2025).
+Experimental observations have been reported in specific platforms, including trapped-ion quantum simulators Ares et al.
+(2023), consistent with theoretical predictions, and single trapped-ion systems Aharony Shapira et al.
 (2024); Zhang et al.
 (2025).
-Motivated by these outstanding works, we want to know whether the quantum Mpemba effect will occur in this open system with diagonal and off-diagonal quasiperiodic modulations as well.
-A previous study has shown that if the initial state is localized, the quantum Mpemba effect will occur Dong et al.
-(2025).
-Therefore, we want to study whether the quantum Mpemba effect will occur if the initial state is not a localized state as well.
-Furthermore, we also aim to provide a unified explanation for the appearance and absence of the quantum Mpemba effect.
+Motivated by recent advances, we examine whether the quantum Mpemba effect can occur in an open, quasiperiodic system with simultaneous diagonal and off-diagonal modulations.
+While earlier work established the effect for localized initial conditions Dong et al.
+(2025), here we consider nonlocalized preparations and delineate the conditions under which the effect, if any, survives.
+We further develop a unified framework that explains both the occurrence and the breakdown of the quantum Mpemba effect.
 This work is organized as follows.
 Section II studies the model with diagonal and off-diagonal modulations and gives out the localization phase diagram.
-Section III studies the thermodynamical applications of the quantum Otto cycle with the quasiperiodic system being the working medium.
-Section IV studies the delocalization-localization transition of the steady state after considering the bond dissipations.
-Section V reveals the phenomenon of quantum Mpemba effect and puts forward a hypothesis to explain the cause of this phenomenon.
+Section III studies the delocalization-localization transition of the steady state after considering the bond dissipations.
+Section IV reveals the phenomenon of quantum Mpemba effect and puts forward a hypothesis to explain the cause of this phenomenon.
 A summary is presented in Sec.
-VI.
+V.
 Figure 1: (Color Online) The phase diagram illustrates how $\log_{10}(\rm MIPR)$ varies with $V/t$ and $t_{1}/t$, where the system size is set to $L=1597$.
@@
 $H=\sum_{n}\left(t_{n}\hat{c}^{{\dagger}}_{n}\hat{c}_{n+1}+H.c.\right)+\sum_{n}V_{n}\hat{c}^{{\dagger}}_{n}\hat{c}_{n},$ (1)
-where $t_{n}=t+t_{1}\cos{(2\pi\alpha n)}$ and $V_{n}=V\cos{(2\pi\alpha n)}$ ($t$ is the unit of energy, $n$ is the site index, and $\alpha=\frac{\sqrt{5}-1}{2}$ is the incommensurate parameter which is the cause of the quasiperiodic modulations.).
+[…]
 Figure 2: (Color Online) (a) $\log_{10}(\rm MIPR)$ as a function of $r$ with $\theta=0$, $\pi/4$, $\pi/3$, and $5\pi/6$.
@@
 The localization phase diagram can be determined from the inverse participation ratio (IPR).
-Given a normalized wave function $\psi_{j}=\sum^{L}_{n=1}\phi_{j}(n)\hat{c}^{{\dagger}}_{n}|0\rangle$ (Here $j$ is the index of wave function), the corresponding ${\rm IPR}_{j}$ is
+Given a normalized wave function $|\psi_{j}\rangle=\sum^{L}_{n=1}\phi_{j}(n)\hat{c}^{{\dagger}}_{n}|0\rangle$ (Here $j$ is the index of wave function), the corresponding ${\rm IPR}_{j}$ is
 ${\rm IPR}_{j}=\sum^{L}_{n=1}|\phi_{j}(n)|^{4}.$ (2)
-It is known that the the $\rm IPR$} of the extended state, critical state, and localized state respectively possess the properties of ${\rm IPR}\approx 0$, $0<{\rm IPR}<1$, and ${\rm IPR}\approx 1$ Biddle and Sarma (2010).
-Under the given parameters, the extended phase, critical phase and localized phase can be further characterized by the average value of $\rm IPR$} of all wave functions, namely the average inverse participation ratio (MIPR), which is denoted as ${\rm MIPR}=\sum^{L}_{j=1}{\rm IPR}_{j}/L$.
-Taking system size $L=1597$, the localization phase diagram in $V$-$t_{1}$ parameter space is plotted in Fig.
+It is known that the $\rm IPR$} of the extended, critical, and localized state, respectively, possess the properties of ${\rm IPR}\approx 0$, $0<{\rm IPR}<1$, and ${\rm IPR}\approx 1$ Biddle and Sarma (2010).
+Under the given parameters, the extended, critical, and localized phase can be further characterized by the average value of $\rm IPR$} of all wave functions, namely the mean inverse participation ratio (MIPR), denoted as ${\rm MIPR}=\sum^{L}_{j=1}{\rm IPR}_{j}/L$.
+With system size $L=1597$, the localization phase diagram in $V$-$t_{1}$ parameter space is plotted in Fig.
 1, where the color bar denotes $\log_{10}(\rm MIPR)$.
-In fact, for the $V=0$ case, the earlier studies have proved that the system exits an extended-critical phase transition Liu et al.
+In fact, for the $V=0$ case, the earlier studies have proved that the system exhibits an extended-critical phase transition Liu et al.
 (2016).
 We can see from Fig.
-1 that under $V=0$, $\log_{10}(\rm MIPR)$ for the extended phase approaching 3-3 and $\log_{10}(\rm MIPR)$ for the critical phase approaches 1-1.
-Based on this, we can infer that in the case of a finite $V$, the quarter elliptic region with $\log_{10}(\rm MIPR)$ close to 3-3 should be the extended phase.
-On its outside, $\log_{10}(\rm MIPR)$ is greater than 1-1, so it should be a localized phase.
-Compared with the non-Hermitian situation Chen et al.
-(2022), the recovery of system’s Hermiticity expands the area of the extended phase.
-To locate the phase boundary of the extended-localized transition, we shall introduce the transformation $r=\sqrt{\frac{V^{2}}{4}+t^{2}_{1}}$ with $\theta={\rm Arctan}(V/t_{1})\in\left[0,2\pi\right]$.
+1(a) that under $V=0$, $\log_{10}(\rm MIPR)$ for the extended phase approaching 3-3 and $\log_{10}(\rm MIPR)$ for the critical phase approaches 1.6-1.6.
+Accordingly, for finite $V$, the quarter-elliptic region in parameter space where $\log_{10}(\rm MIPR)\approx-3$ is identified with the extended (delocalized) phase, whereas outside this region $\log_{10}(\rm MIPR)>-1$ indicates localization.
+To locate the phase boundary of the extended-localized transition, we shall introduce the transformation $r=\sqrt{V^{2}/4+t^{2}_{1}}$ and $\theta={\rm Arctan}(V/t_{1})$ ($\in\left[0,2\pi\right]$).
 Under $L=1597$, the $\log_{10}(\rm MIPR)$ versus $r$ with different $\theta$ are plotted in Fig.
 2(a).
-Intuitively, we can see that for different $\theta$, the $\log_{10}(\rm MIPR)$-$r$ curves characterizing the extended-localized (critical) transition all present a jump at the critical value $r_{c}/t=1$, namely the phase boundary $\sqrt{\frac{V^{2}}{4}+t^{2}_{1}}=t$, which has been plotted as the red dashed line in Fig.
+Intuitively, we can see that for different $\theta$, the $\log_{10}(\rm MIPR)$-$r$ curves characterizing the extended-localized (critical) transition all present a jump at the critical value $r_{c}/t=1$, namely the phase boundary $\sqrt{V^{2}/4+t^{2}_{1}}=t$ extracted from the numerical results, which has been plotted as the red dashed line in Fig.
 1.
-To further verify the above mentioned conclusions we have drawn, we introduce the fraction dimension $D$.
-Consider a system where the system size $L$ equals the $m$-th Fibnacci number $F_{m}$ and the incommensurate parameter $\alpha$ is replaced by $\alpha=F_{m-1}/F_{m}$.
+To further verify the above mentioned conclusions, we introduce the fraction dimension $D$.
+Consider a system where the system size $L$ equals the $m$-th Fibonacci number $F_{m}$ and the incommensurate parameter $\alpha$ is replaced by $\alpha=F_{m-1}/F_{m}$.
 At a specific lattice site $n$, the fraction dimension $D_{n}$ can be derived from the formula
@@
 For a critical state, $D_{n}$ lies within a finite interval $\left[D_{min},D_{max}\right]$.
-Thus, the minimum value $D_{min}$ directly reflects the nature of a given wave function $\psi_{j}$.
-Specifically, $D_{min}\rightarrow 0$ indicates a localized state, $0<D_{min}<$ points to a critical state, and $D_{min}\rightarrow 1$ signifies an extended state.
+Thus, the minimum value $D_{min}$ directly reflects the nature of a given wave function $|\psi_{j}\rangle$.
+Specifically, $D_{min}\rightarrow 0$ indicates a localized state, $0<D_{min}<1$ points to a critical state, and $D_{min}\rightarrow 1$ signifies an extended state.
 For generality, we use the average of $D_{min}$ over all states, defined as $\overline{D_{min}}=\sum^{L}_{j=1}D^{j}_{min}/L$.
@@
 2(b), we observe that the relevant $D_{min}$ approaches $1$ at the parameter sites $\left(r,\theta\right)=\left(0.5t,\frac{\pi}{4}\right)$ and $\left(r,\theta\right)=\left(0.8t,\frac{\pi}{3}\right)$.
-This observations confirm that the system resides in the extended phase under these conditions.
+These observations confirm that the system resides in the extended phase under these conditions.
 As anticipated, the corresponding $D_{min}$ values fall within the interval $\left(0,1\right)$ in the thermodynamic limit when $\left(r,\theta\right)=\left(1.5t,0\right)$ and $\left(r,\theta\right)=\left(2t,0\right)$.
-These results clearly reveal the systems’s critical-phase characteristics in such parameter regimes.
+These results clearly reveal the critical-phase characteristics of the system in such parameter regimes.
 When $\left(r,\theta\right)=\left(1.5t,\frac{\pi}{3}\right)$ and $\left(r,\theta\right)=\left(2t,\frac{\pi}{6}\right)$, the associated $D_{min}$ extrapolates to 0.
 This outcome identifies that the system is in the localized phase at these parameter points.
-III Thermodynamical applications
-In recent years, some researches have been carried out on the thermodynamical applications of quasi-periodic systems Chiaracane et al.
-(2020); Xu and Cheng (2025), leading to the identification of their applications in aspects such as quantum heat engine, quantum heater, and quantum accelerator.
-Specifically, it has been demonstrated that the extended phase plays a role in sustaining the operational mode of heat engine, whereas the critical phase is conducive to maintaining the working modes of quantum heater and quantum accelerator.
-Within this section, we focus on investigating the thermodynamic applications of this generalized Aubry-André (AA) model with both diagonal and off-diagonal quasiperiodic modulations.
-Our primary objectives include verifying whether the extended phase of this model still contributes to preserving the working mode of quantum heat engines, as well as determining which specific working mode the localized phase favors.
-Figure 3: (Color Online) Schematic illustration of the four-stroke quantum heat cycle process.
-$T_{h}$ and $T_{c}$ stand for the high-temperature and low-temperature heat sources, respectively.
-The working medium employed is the generalized AA model with diagonal and off-diagonal quasiperiodic modulations.
-$r^{i}$ and $r^{f}$ represent the systematic parameters of the corresponding Hamiltonians.
-The system size is $L=610$.
-We leverage this quasiperiodic system as the working medium to construct a quantum heat cycle process.
-The schematic of this cycle is shown in Fig.
-3.
-This quantum heat cycle can be regarded as the quantum counterpart Quan et al.
-(2007); Niskanen et al.
-(2007); Abah et al.
-(2012); Deng et al.
-(2013); del Campo et al.
-(2014); Roßnagel et al.
-(2016); Zheng et al.
-(2016); Kosloff and Rezek (2017); Deffner and Campbell (2019); Campisi and Fazio (2016); Buffoni et al.
-(2019); Solfanelli et al.
-(2020); Zhang et al.
-(2022) of the classical Otto cycle.
-The first (④ $\rightarrow$ ①) and third (② $\rightarrow$ ③) strokes occur under thermal contact with high-temperature ($T_{c}$) heat reservoirs.
-These strokes proceed without external driving or particle exchange.
-The second (① $\rightarrow$ ②) and fourth (③ $\rightarrow$ ④) strokes are adiabatic from a thermodynamic viewpoint (meaning thermal isolation), yet they might not be strictly adiabatic in the quantum-mechanical sense, since quantum transitions can occur during the evolution [82].
-In the first stroke, the working medium, characterized by the Hamiltonian $H(r^{i})$, will eventually reach thermal equilibrium and relax into a Gibbs state.
-The density matrix $\rho_{1}$ of this state is given by $\rho_{1}=\frac{e^{-\beta_{h}H(r^{i})}}{Z_{1}}$, where $\beta_{h}=\frac{1}{k_{b}T_{h}}$ ($k_{b}$ is Boltzmann constant) and $Z_{1}={\rm Tr}\left[e^{-\beta_{h}H(r^{i})}\right]$ is the partition function.
-Thus, at thermal equilibrium, the system energy is $E_{1}={\rm Tr}\left[\rho_{1}H(r^{i})\right]$.
-During the second stroke, the Hamiltonian parameter is switched from $r^{i}$ to $r_{f}$.
-In this stroke, only work is done, and no heat exchange takes place.
-The Gibbs state $\rho_{2}$ of this stroke remains unchanged, i.e, $\rho_{2}=\rho_{1}$, but the energy $E_{2}$ becomes $E_{2}={\rm Tr}\left[\rho_{2}H(r)\right]$.
-In the third stroke, the medium with $r=r^{f}$ comes into contact with a heat source having $\beta_{c}=\frac{1}{k_{b}T_{c}}$.
-As a result, the Gibbs state $\rho_{3}=\frac{e^{-\beta_{c}H(r^{f})}}{Z_{2}}$, with the partition function $Z_{2}={\rm Tr}\left[e^{-\beta_{c}H(r^{f})}\right]$.
-Then, the energy of the medium is $E_{3}={\rm Tr}\left[\rho_{3}H(r^{f})\right]$.
-The fourth stroke can be regarded as the thermal annealing process.
-In this process, the hopping parameter in the medium is changed back from $r^{f}$ to $r^{i}$.
-However, the Gibbs state $\rho_{4}$ stays as $\rho_{4}=\rho_{3}$.
-Hence, the energy of the medium becomes $E_{4}={\rm Tr}\left[\rho_{4}H(r^{i})\right]$.
-Figure 4: (Color Online) Working modes of the four-stroke cycle.
-(a) $r^{f}=0.5t$ and $\theta=\pi/6$.
-(b) $r^{f}=0.5t$ and $\theta=\pi/4$.
-(c) $r^{f}=0.9t$ and $\theta=\pi/6$.
-(d) $r^{f}=0.9t$ and $\theta=\pi/4$.
-(e) $r^{f}=1.1t$ and $\theta=\pi/6$.
-(f) $r^{f}=1.1t$ and $\theta=\pi/4$.
-The blue regions represent the Heat engine.
-The yellow regions denote the Accelerator and the brown regions denote the Heater.
-Other parameters are $L=610$ and $k_{b}T_{c}=0.1t$.
-After the working medium undergoes this complete cycle, we can determine the heat absorbed $Q_{h}$ ($Q_{h}=E_{1}-E_{4}$) from the $T_{h}$ source, the heat released $Q_{c}$ ($Q_{c}=E_{3}-E_{2}$) to the $T_{c}$ source, and the net work $W$ done by the working medium, where $W=Q_{h}+Q_{c}$.
-Importantly, the heat cycle process abides by the Clausius inequality, a cornerstone of thermodynamics.
-Depending on the values of $Q_{h}$, $Q_{c}$, and $W$, the engine using the extended-localized (critical) quasiperiodic system as the working medium exhibits different modes Buffoni et al.
-(2019); Solfanelli et al.
-(2020):
-  1. 1.
-    Heat engine: $Q_{h}>0$, $Q_{c}<0$, and $W>0$;
-  2. 2.
-    Refrigerator: $Q_{h}<0$, $Q_{c}>0$, and $W<0$;
-  3. 3.
-    Heater: $Q_{h}<0$, $Q_{c}<0$, and $W<0$;
-  4. 4.
-    Accelerator: $Q_{h}>0$, $Q_{c}<0$, and $W<0$.
-For a system with size $L=610$ and $k_{b}T_{c}=0.1t$, after analyzing the values of $Q_{h}$, $Q_{c}$ and $W$, the corresponding working modes of the four-stroke cycle under different $r^{f}$ and $\theta$ are plotted in Figs.
-4(a)-4(f), respectively.
-It can be observed that this cycle exhibits a diverse range of working modes.
-The blue regions represent the Heat engine mode.
-The yellow regions correspond to the Accelerator mode, while the brown regions indicate the Heater working mode.
-It is noted from Figs.
-4(a)-4(d) that when $r^{f}$ and $r^{i}$ are below the critical value $r_{c}$, the Heat engine mode appears.
-This suggests that similar to the results of the critical-extended quasi-periodic system Xu and Cheng (2025), for the quasi-periodic system under a finite $V$, its extended phase is favorable for the design of a quantum heat engine as well.
-As $r^{f}$ increases and surpasses the critical values $r_{c}=t$, the system entries the localized phase.
-From Figs.
-4(e) and 4(f), we can see that regions representing the Heater mode expand.
-This indicates that the localized phase is more conducive to the realization of a quantum heater.
-Additionally, it is clear that there are extensive parameter ranges corresponding to the Accelerator mode, regardless of whether the system is in the extended phase or the localized phase.
-Combining with previous research on extended-critical quasiperiodic system Xu and Cheng (2025), we draw a conclusion that all the extended, critical and localized phases facilitate the realization of a quantum accelerator, and both the localized and critical phases are conducive to the realization of quantum heater.
-Moreover, the results shown in Fig.
-4 offer strategies for regulating the working mode of the four-stroke cycle.
-The transition between different working modes can be achieved by tuning $r^{f}$, $r^{i}$, and $k_{b}T_{h}$.
-For example, when $r^{f}$ is much smaller than $r_{c}$ ( refer to Figs.
-4(a) and 4(b)), or vice versa, there are two working modes, i.e., Accelerator and Heat engine.
-By tuning $r^{i}$ and $k_{b}T_{h}$, the Otto cycle can switch between the two working modes.
-When $r^{f}$ is close to but still less than $r^{c}$ and $k_{b}T_{h}$ is small (refer to Figs.
-4(c) and 4(d)), there are three working modes.
-Thus, the cycle’s working mode will change from Heater to Accelerator and then to Heat engine when we gradually increase $r^{i}$.
-When $r^{f}$ is larger than $r_{c}$, the cycle can be toggled between Heater and Accelerator modes by tuning $r^{i}$ or $k_{b}T_{h}$.
-IV Dissipation driven delocalization-localization transition
-In the preceding section, we conducted a systematic investigation into the thermodynamic characteristics of this quasi-periodic system under thermal equilibrium conditions.
-Our findings revealed two key insights: firstly, the system’s thermodynamic behaviors are tied to its initial properties; secondly, the system holds potential application value for developing quantum heat engine, heater, and accelerator.
-We also analyzed approaches to regulate these three working modes.
-In addition to the thermodynamic aspect, recent research efforts have yielded conflicting perspectives regarding the steady-state delocalization-localization transition under dissipative modulation Longhi (2024a); Liu et al.
+III Dissipation driven delocalization-localization transition
+Recent research efforts have yielded conflicting perspectives regarding the steady-state delocalization-localization transition under dissipative modulation Longhi (2024); Liu et al.
 (2024); Peng et al.
-(2024b); Hu et al.
-(2025b).
+(2024) .
 By employing the dissipation, one can precisely control the delocalization and localization properties of the system under non-equilibrium steady states.
 Therefore, it motivates us to study the delocalization and localization properties of the steady state after introducing the dissipation.
-Besides, some studies suggest this transition is independent of the system’s initial properties.
+Besides, some studies suggest this transition is independent of the system’s initial properties Liu et al.
+(2024).
 Alternative viewpoints argue that the system’s initial attributes can influence the transition.
@@
 et al.
-(2017); Yusipov et al.
-(2018); Vakulchyk et al.
 (2017); Jiang et al.
@@
 $\dot{\rho}=-\left[H,\rho\right]+\sum_{n}\gamma_{n}\mathcal{D}\left[L_{n}\right]\rho,$ (4)
-Here, the dissipator $\mathcal{D}\left[L_{n}\right]$ is defined as $\mathcal{D}\left[L_{n}\right]=L_{n}\rho L^{{\dagger}}_{n}-\{L^{{\dagger}}_{n}L_{n},\rho\}/2$, and $\gamma_{n}=\gamma\cos({2\pi\alpha n})$ represents the tunably quasiperidoic bond dissipation strength with the incommensurate parameter $\alpha=\frac{\sqrt{5}-1}{2}$.
-The rationale for our selection of the quasiperiodic form for the key dissipation operator stems from the consideration of the following aspect: the initial key dissipation operator is made up of four terms, with two being dephasing terms Longhi (2024a) and the remaining two being hopping terms.
+Here, the dissipator $\mathcal{D}\left[L_{n}\right]$ is defined as $\mathcal{D}\left[L_{n}\right]=L_{n}\rho L^{{\dagger}}_{n}-\{L^{{\dagger}}_{n}L_{n},\rho\}/2$, and $\gamma_{n}=\gamma\cos({2\pi\alpha n})$ represents the tunable quasiperiodic bond dissipation strength with the incommensurate parameter $\alpha=\frac{\sqrt{5}-1}{2}$.
+The rationale for our selection of the quasiperiodic form for the key dissipation operator stems from the consideration of the following aspect: the initial key dissipation operator is made up of four terms, with two being dephasing terms Longhi (2024) and the remaining two being hopping terms.
 As indicated in Ref.
@@
 Moreover, this bond dissipation operator can be experimentally realized by introducing the auxiliary lattice Diehl et al.
-(2008); Kraus et al.
-(2008); Diehl et al.
-(2010, 2011); Bardyn et al.
-(2013).
+(2008, 2010, 2011).
 Owing to these factors, we utilized such a dissipative operator.
@@
 With $L=100$ and $\ell=1$, the delocalization phase diagrams for $V=0$, $V=0.5t$, $V=t$, and $V=2t$ are plotted in Figs.
-5(a)- 5(d), respectively.
+3(a)- 3(d), respectively.
 The phase diagram is obtained by analyzing the position of the centroid in a steady state.
-We define a steady state where the centroid position is no more than five times the lattice constant from the system boundary as a localized state; otherwise, it is a delocalized state.
-The reason for choosing such a five-fold distance as the criterion is based on the experimental fact that in Roati et al.
-(2008), the authors reproduced a one-dimensional AA model in the cold atom optical lattice experimental system and observed the Anderson localization phase transition, where the wave packet broadening of the localized state was approximately $10$ times the lattice constant Roati et al.
-(2008).
+We define a steady state where the centroid position (its definition is given in Eq.(9)) is no more than five times the lattice constant from the system boundary as a localized state; otherwise, it is a delocalized state.
+Our choice of criterion is guided by the experiment reported in Ref.
+Roati et al.
+(2008), which realized a one-dimensional Aubry–André (AA) model in a cold-atom optical lattice and observed the Anderson-localization transition; in the localized regime, the wave packet of an initially localized state broadened by approximately 10 lattice constants.
 Therefore, it is feasible to distinguish the delocalization-localization transition of the steady state through the position of the center of mass.
 As can be seen from Figs.
-5(a)-5(c), when $t_{1}$ is less than $0.5t$, there is a large parameter range that makes all the steady states delocalized.
+3(a)-3(c), when $t_{1}$ is less than $0.5t$, there is a large parameter range that makes all the steady states delocalized.
 This indicates that under these parameters, the steady state can retain the initial delocalized property of the system.
@@
 It is indicated that adjusting the parameter $V$ can also regulate the localization property of the steady state.
-For instance, when V is at a relatively small value, it can be seen from 5(a) and 5(b) that when $t_{1}>t$, there are still delocalized states within it, resulting in the reentrance delocalization phenomenon.
+For instance, when V is at a relatively small value, it can be seen from 3(a) and 3(b) that when $t_{1}>t$, there are still delocalized states within it, resulting in the reentrance delocalization phenomenon.
 Moreover, the value of $V$ is smaller, making the delocalization region larger.
 When $V$ is relatively large, it can be seen from Figs.
-5(c) and 5(d) that the delocalized region disappears.
+3(c) and 3(d) that the delocalized region disappears.
 In particular, we find that when $V=2t$, even if the initial system is in the localized phase, the steady state is still delocalized (as can be seen from Fig.
-5(d), the delocalized state exists within a small range of $t_{1}$ and $\gamma$).
-V quantum Mpemba effect and starting-line hypothesis
-Regarding to the delocalization-localization transition in open quasiperiodic system, the focus on this topic mainly lies in the localization properties of steady states.
+3(d), the delocalized state exists within a small range of $t_{1}$ and $\gamma$).
+IV quantum Mpemba effect and starting-line hypothesis
+Regarding the delocalization-localization transition in open quasiperiodic system, the focus on this topic mainly lies in the localization properties of steady states.
 For an anomalous dynamic process: the relaxation rate of a system whose initial state is further from the equilibrium state is actually faster than that of a system whose initial state is closer to the equilibrium state, that is, the quantum Mpemba effect, remains to be studied and discussed.
@@
 (d) $V=0.2t$, $t_{1}=1.5t$, and $\gamma=0.2t$.
-(e) $V=0.2t$, $t_{1}=0.8$, $\gamma=0.2t$.
-(f) $V=0.5t$, $t_{1}=0.5$, $\gamma=0.2t$.
-The blue regions denote the delocalization phases.
-The yellow regions denote the localization phases.
+(e) $V=0.2t$, $t_{1}=0.8t$, $\gamma=0.2t$.
+(f) $V=0.5t$, $t_{1}=0.5t$, $\gamma=0.2t$.
+The ordinal number in the caption indicates the ordinal number of the initial state in the eigenstates of $H$.
 Other parameter is $L=100$.
-We employ the density matrix $\rho(\tau)$ that describe the evolution of the system over time and the steady state density matrix $\rho_{ss}$ to calculate the standard distance between the matrices, this is, the trace distance, to characterize the quantum Mpemba effect.
+We employ the density matrix $\rho(\tau)$ that describes the evolution of the system over time and the steady state density matrix $\rho_{ss}$ to calculate the standard distance between the matrices, this is, the trace distance, to characterize the quantum Mpemba effect.
 The trace distance $d_{\rm Tr}$ is defined as
@@
 where $T=\rho(\tau)-\rho_{ss}$.
-Under Lindbladian dynamics, this quantity displays monotonic behavior over time.
-Notably, it has already been utilized as an experimental probe to investigate both the strong quantum Mpemba effect and its inverse counterpart Aharony Shapira et al.
+Under Lindbladian dynamics, this quantity is monotonic in time and has been employed experimentally to probe both the strong quantum Mpemba effect and its inverse counterpart Aharony Shapira et al.
 (2024); Zhang et al.
 (2025).
-Prior to investigating the quantum Mpemba effect, we first introduce the ‘starting-line hypothesis’.
-This hypothesis posits the following: for an initial state whose energy is closest to the steady-state energy, if the distance between its center of mass and the steady-state center of mass is larger than that of other initial states with energies further from the steady-state energy, then the quantum Mpemba effect will arise.
-Before verifying this hypothesis, we can draw a simple inference regarding which initial states are capable of triggering the quantum Mpemba effect.
-For quantum states with extended properties, the position of their center of mass is consistently situated at the center of the system.
-In contrast, for quantum states with localized and critical properties, the positions of their centroids are not fixed at the center of system; instead, they can be distributed at other locations within the system.
-As a result, when quantum states with localized and critical properties are chosen as initial states, they will encounter the ‘starting-line discrepancy’ a condition that gives rise to the quantum Mpemba effect.
-On the other hand, for initial states with extended properties, their centroids all align at the same ‘starting-line’ (i.e., the center of system), meaning the quantum Mpemba effect will not occur.
-We take three different types of system eigenstates as the initial states and calculated the evolution of trace distance over time, as shown in Figures 6(a)-6(f).
+A nonzero trace distance $d_{\rm{Tr}}>0$ indicates that the system has not yet relaxed to the steady state, whereas $d_{\rm{Tr}}=0$ certifies convergence to the steady state.
+Before examining the quantum Mpemba effect, we introduce the starting-line hypothesis.
+Consider a set of initial states characterized by their energies and center-of-mass (COM) positions.
+Let $E_{\mathrm{ss}}$ and $x_{\mathrm{com}}^{\mathrm{ss}}$ denote the energy and COM of the steady state, and define the COM distance as
+$\Delta x_{\mathrm{com}}\equiv\left|x_{\mathrm{com}}(0)-x_{\mathrm{com}}^{\mathrm{ss}}\right|,$ (7)
+[…]
+[…]
+[…]
+[…]
+where $n$s represent as the marks of the site index.
+The hypothesis posits: if the state whose energy is close to $E_{\mathrm{ss}}$ has a larger $\Delta x_{\mathrm{com}}$ than another state whose energy lies farther from $E_{\mathrm{ss}}$, then the latter can relax faster, manifesting the quantum Mpemba effect.
+This yields a simple criterion for which initial states can trigger the effect.
+For extended states, $x_{\mathrm{com}}(0)$ is pinned to the system center, so candidate extended initial states share almost the same ’starting line’ and $\Delta x_{\mathrm{com}}$ do not exhibit the effect.
+By contrast, localized and critical states generally have $x_{\mathrm{com}}(0)$ displaced from the center; such ’starting-line discrepancies’ enable a state that is farther in energy (but closer in COM) to relax faster than a state that is closer in energy (but farther in COM), thereby producing the quantum Mpemba effect.
+We take three different types of system eigenstates as the initial states and calculated the evolution of trace distance over time, as shown in Figures 4(a)-4(f).
 The ordinal number in the caption indicates the ordinal number of the initial state in the eigenstates of $H$.
-Specifically, the initial state corresponding to the red curve is the eigenstate of the system that is closest to the steady-state energy.
-Figures 6(a) and 6(b) show the results of the local state as the initial state.
-The centroid positions corresponding to the initial state (rounded off, the same below) are respectively: $n=\{60,49,39\}$ and $n=\{60,49,41\}$.
-Figures 6(c) and 6(d) show the results of delocalized states with critical properties as the initial states, where the corresponding centroid positions of the initial states are respectively: $n=\{63,61,30,83,49\}$ and $n=\{66,63,15,59,71\}$.
-Figures 6(e) and 6(f) present delocalized states with critical properties as the results of initial states, where the corresponding centroid positions of the initial states are respectively: all $n=50$ and all $n=51$.
-We can observe that for the initial states with extended properties, the dynamical evolution behavior does not exhibit the quantum Mpemba effect, while for the initial states with critical and localized properties, the dynamical evolution behavior will exhibit the quantum Mpemba effect.
-From the perspective of the centroid position of the initial state, when the Mpemba effect occurs, the wave packet of the initial state is relatively closer to the boundary of the system, which confirms the starting-line hypothesis we proposed.
-VI Summary
-For one-dimensional chain models where both hoppings and onsite potential are quasiperiodically modulated, we have analyzed their delocalization-localization transitions under equilibrium and non-equilibrium conditions, thermodynamic applications, and dynamic evolution behaviors.
-We numerically calculated and obtained the delocalization-localization phase diagram of the model, and verified it through MIPR and fractal dimensions.
-The results show that the delocalized phase with extended feature and and the localized phase can be separated by an ellipse curve.
-By using the model as the working medium of the quantum Otto cycle, we analyzed the underlying working modes of the cycle under multiple sets of typical parameters.
-The results show that when the system is in the delocalized phase with extended feature, it is conducive to the realization of the quantum heat engine, while being in the localized phase is conducive to the realization of the quantum heater.
-Furthermore, regardless of the phase region where the system is located, the working parameter space of the accelerator mode is always the largest, while that of the refrigerator mode never exists.
-After calculating the delocalization-localization phase diagram of the non-equilibrium steady state, we discussed the influence of quasiperiodic bond dissipation on the initial state evolution of the model and found that under moderate $V$ and $t_{1}$, the steady state of the system can largely maintain the initial delocalization property of the system.
-By adjusting $\gamma$ and $t_{1}$ or by adjusting the parameter $V$, the properties of the steady state can be regulated, gradually changing it from a delocalized state to a localized state.
-When the value of $V$ is much smaller, even if the initial state is localized, the steady state can also be delocalized.
-We analyzed the non-equilibrium dynamic evolution behavior and the existence of the quantum Mpemba effect using the trace distance.
-The results told that the quantum states with localized and critical properties as initial states are necessary conditions for the occurrence of the quantum Mpemba effect.
-These results can be explained by the starting-line hypothesis we have proposed.
-As for whether this hypothesis still holds true for other open quasi-periodic systems, this is an open question that remains to be tested.
+Specifically, the initial state energy corresponding to the red curve is closer to the steady-state energy than that of other initial states.
+Figures 4(a) and 4(b) show the results of the delocalized states with critical properties as the initial state.
+The COM positions corresponding to the initial states (rounded off, the same below) are respectively: $n=\{60,49,39\}$ and $n=\{60,49,41\}$, and the COM distances are respectively: $\Delta x_{\mathrm{com}}=\{57,47,36\}$ and $\Delta x_{\mathrm{com}}=\{56,46,38\}$.
+It can be seen from the data that the COM distance $\Delta x_{\mathrm{com}}$ between the initial state with energy close to the steady state and the steady state is actually greater than that when the energy is far from the steady state.
+The time required for $d_{\rm Tr}$ to evolve to zero is longer, and thus the so-called quantum Mpemba effect emerges.
+Figures 4(c) and 4(d) show the results of the localized states as the initial states, where the corresponding COM positions of the initial states are respectively: $n=\{63,61,30,83,49\}$ and $n=\{66,63,15,56,71\}$, and the COM distances are respectively: $\Delta x_{\mathrm{com}}=\{60,58,27,80,46\}$ and $\Delta x_{\mathrm{com}}=\{61,58,10,51,66\}$.
+We can observe that when energy moves away from the initial state of steady-state energy, it evolves to the steady state more quickly compared to the quantum state where energy is close to the initial state energy.
+Therefore, the so-called quantum Mpemba effect also emerges.
+We can still observe that for the initial state where energy is far from the steady-state energy, it evolves to the steady state more quickly than the quantum state where energy is close to the steady-state energy.
+Therefore, the so-called quantum Mpemba effect emerges.
+This phenomenon can be explained by the proposed hypothesis as well, because the COM of the initial state with energy close to the steady state is further away from the steady state compared to other initial states, and thus it takes longer time to evolve to the steady state.
+Figures 4(e) and 4(f) present delocalized states with extended properties as the results of initial states, where the corresponding COM positions of the initial states are respectively: all $n=50$ and all $n=51$, and the COM distances are respectively: $\Delta x_{\mathrm{com}}=42$ and $\Delta x_{\mathrm{com}}=21$.
+We can see that regardless of whether the energy is close to or far from the initial state of the steady-state energy, the time it takes for $d_{\rm Tr}$ to evolve to zero is almost indistinguishable, and thus the so-called quantum Mpemba effect does not occur.
+This phenomenon can be explained by using our hypothesis.
+Because the initial COM positions and COM distances of different initial states are almost the same, the time required for evolution to a steady state is also almost the same.
+Based on these results, we can conclude that for the initial states with extended properties, the dynamical evolution behavior does not exhibit the quantum Mpemba effect, while for the initial states with critical and localized properties, the dynamical evolution behavior will exhibit the quantum Mpemba effect.
+From the perspective of the COM position of the initial state, when the Mpemba effect occurs, the wave packet of the initial state is relatively closer to the boundary of the system, which confirms the starting-line hypothesis we proposed.
+V Summary
+We have investigated a one-dimensional tight-binding model with simultaneous quasiperiodic modulations of both hopping and onsite potential, focusing on localization–delocalization transitions under equilibrium and nonequilibrium conditions, and quantum relaxation dynamics.
+Using finite-size numerics, we have mapped the equilibrium phase diagram and verified it via the mean inverse participation ratio and fractal dimensions, finding that the delocalized (extended) and localized phases are separated by an approximately elliptical phase boundary.
+By computing the nonequilibrium steady-state phase diagram under quasiperiodic bond dissipation and analyzing initial-state dependence, we have known: for moderate $V$ and $t_{1}$, the steady state largely inherits the initial state’s delocalized character; tuning $\gamma$, $t_{1}$ or $V$ continuously drives the steady state from delocalized to localized, and for sufficiently small $V$ the steady state can become delocalized even from localized initial conditions.
+Finally, using the trace distance, we have demonstrated that there exists quantum Mpemba effect in which states initialized in localized or critical regimes can relax faster than states prepared closer to the steady state.
+These observations are rationalized by a starting-line hypothesis, whose applicability to broader classes of open quasiperiodic systems remains an open question.
 This research is supported by Zhejiang Provincial Natural Science Foundation of China under Grant No.